OpenStudy (anonymous):

Solve in C : z^6 = -1

jimthompson5910 (jim_thompson5910):

Hint: z^6 = -1 z^6 + 1 = 0 (z^2)^3 + 1^3 = 0 (z^2 + 1)((z^2)^2 + (z^2)*1 + 1^2) = 0 (z^2 + 1)(z^4 + z^2 + 1) = 0

OpenStudy (anonymous):

thank you z=i or z=-i but (z^4 + z^2 + 1) = 0 let z^2 =t t in C t^2-t+1=0 delta =......

jimthompson5910 (jim_thompson5910):

use the quad formula to solve for t in t^2-t+1=0

jimthompson5910 (jim_thompson5910):

oh made a typo, should be z^4 - z^2 + 1

jimthompson5910 (jim_thompson5910):

but you caught it

OpenStudy (anonymous):

delta = b^2-4ac and a =1 b=-1 c=1

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

delta = -3=3i^2

jimthompson5910 (jim_thompson5910):

not sure what you mean there

OpenStudy (anonymous):

I am sure